OpenStudy (anonymous):

How do I simplify: sqr. root 3 / 4 sqr. root 10?

OpenStudy (anonymous):

\[\sqrt{3} \div 4\sqrt{10}\]

OpenStudy (anonymous):

multiply numerator and denominator by sqrt(10), then you will have rationalized the denominator

OpenStudy (anonymous):

I got \[\sqrt{30} / 40\]

OpenStudy (anonymous):

thats the right answer

OpenStudy (anonymous):

but the back of my text book says: \[\sqrt{30} / 4\]

OpenStudy (anonymous):

your book is wrong

OpenStudy (anonymous):

unless the question is 10sqrt(3)/(4sqrt(10))

OpenStudy (accessdenied):

i found a typo in one of my books previously (calculus book had one of the intervals backwards), its not impossible that they made a mistake!

OpenStudy (anonymous):

yes and they piss you off when you keep going back over your work to see where you went wrong

OpenStudy (anonymous):

ohhh, the question I posted is correct & thank you!